OpenStudy (anonymous):

Hi everyone! I need help with this question. It might take me awhile to write it so please be patient! thanks again if your helping me!

OpenStudy (anonymous):

\[\left| m \right| +2 =11\] When i solve this I get ( 9,-13) but on my answer sheet it says the answer is (9-9)

OpenStudy (anonymous):

@lopezking1

OpenStudy (anonymous):

@Preetha @IMStuck

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

your first step is to subtract 2 from both sides

jimthompson5910 (jim_thompson5910):

that will give you |m| = 9 that breaks down into m = 9 or m = -9 since both numbers are nine units away from 0

jimthompson5910 (jim_thompson5910):

you would NOT do this |m|+2 = 11 m+2 = 11 or m+2 = -11 m = 11-2 or m = -11-2 m = 9 or m = -13 since you need to subtract 2 first, then break up the absolute value. You would only follow those steps if you had |m+2| = 11

OpenStudy (anonymous):

thanks I see my mistake.!!!

OpenStudy (anonymous):

hi